OpenStudy (anonymous):

√-4(6+√-25) multiply, answer in a+bi form.

OpenStudy (lalaly):

2i(6+5i) = 12i-10

OpenStudy (lalaly):

a=-10 b=12

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{-4}(6+\sqrt{-25})\] \[\large \sqrt{-1*2^2}(6+\sqrt{-1*5^2})\] \[\large \sqrt{-1}*\sqrt{2^2}(6+\sqrt{-1}*\sqrt{5^2})\] \[\large i*2(6+i*5)\] \[\large 2i(6+5i)\] \[\large 2i(6)+2i(5i)\] \[\large 12i+10i^2\] \[\large 12i+10(-1)\] \[\large 12i-10\] \[\large -10+12i\] So \[\large \sqrt{-4}(6+\sqrt{-25})=-10+12i\]